OpenStudy (anonymous):

Pretty dumb to ask, but how do you divide: x^4 +5x^3 -2x^2 +5x -3 by x^2 +1?

OpenStudy (anonymous):

HOW do you divide it is what i'm asking..

OpenStudy (anonymous):

\[\frac{ x^4 + 5x^3 - 2x^2 + 5x - 3 }{ x^2 + 1 }\]

OpenStudy (anonymous):

....yes. I know that. what would the simplified form be...

OpenStudy (anonymous):

dumber ans would be --> you simple dont! ;)

OpenStudy (anonymous):

and the smarter thing would be ,, long division..

OpenStudy (lgbasallote):

are you familiar with polynomial division @breberea ?

OpenStudy (anonymous):

Not at all. I normally do synthetic, but I don't know how to do that with x^2+1..

OpenStudy (lgbasallote):

you make a point...

OpenStudy (lgbasallote):

like @Aryang said...you simply dont

OpenStudy (lgbasallote):

synthetic can only be done with linear dividends...in this case you have quadratic...so you can't use synthetic

OpenStudy (lgbasallote):

you should just use long division

OpenStudy (anonymous):

@lgbasallote that's the thing. I don't know how to do long division with equations like this. That's why I said it's pretty dumb..

OpenStudy (anonymous):

if you want to avoid lengthy work,,you may simply numerator,, for example in x^4 +5x^3 -2x^2 +5x -3 we can write it as x^4 + x^2 + 5x^3 - 3x^2 + 5x -3 now first 2 terms here are divisible by x^2 +1 rest you have 5x^3 - 3x^2 + 5x -3 = 5x^3 + 5x - 3x^2 - 3 again first 2 terms are divisible by x^2 + 1 rest remaining is -3x^2 - 3 = -3(x^2 +1) hope that helped..

OpenStudy (anonymous):

@Aryang I'm sure you're right, but I cant do it like that. It's either: f(x) = (x2 + 1)( x2 + 5x - 3) f(x) = (x2 + 1)( x2 + 5x - 3) - 15x - 9 f(x) = (x2 + 1)( x2 - 5x - 3) - 15x - 9 f(x) = (x2 + 1)( x2 - 5x - 3)